# Filling in the Blanks:

### v1

@@ Appendix D Phase-Clumped or Phase-Mixed
Figure 19: The cylindrical velocity distributions of the different stellar Halo groups, as recovered by our methodology (lefthand side), against the true distributions (righthand side). Different groups are represented using different colours, with the opacity of the colours indicating $\left(0.5\sigma,1\sigma,2\sigma\right)$ levels (equivalently $38\%,64\%,95\%$). This figure is a comparison to Fig. 10, where the phase-clumped nature of the selected groups are additionally modelled with an additional step. Here are the results of assuming all groups are perfectly phase-mixed, resulting in near-symmetric velocity distributions in the radial and vertical components.

### v2

@@ Appendix D Phase-clumped or phase-mixed
Figure 19: Cylindrical velocity distributions of the different stellar Halo groups, as recovered by our methodology (lefthand side), against the true distributions (righthand side). Different groups are represented using different colours, with the opacity of the colours indicating $\left(0.5\sigma,1\sigma,2\sigma\right)$ levels (equivalently $38\%,64\%,95\%$). This figure is a comparison to Fig. 10, where the phase-clumped nature of the selected groups are additionally modelled with an additional step. Here are the results of assuming all groups are perfectly phase-mixed, resulting in near-symmetric velocity distributions in the radial and vertical components.